OpenStudy (anonymous):

What is the equation of the lne that passes through points (5,4) and (-2,0)?

OpenStudy (anonymous):

rise/run

OpenStudy (anonymous):

first find slope

OpenStudy (anonymous):

then use y=mx+b

OpenStudy (anonymous):

subsitute slope

OpenStudy (anonymous):

and then find y intercept

OpenStudy (mathstudent55):

Call point (5, 4) = (x1, y1), and call point (-2, 0) = (x2, y2) Slope = m = (y2 - y1)/(x2 - x1) The equation of the line is: y - y1 = m(x - x1)

OpenStudy (anonymous):

he has point slope form

OpenStudy (anonymous):

both can be used though

OpenStudy (anonymous):

ok. thank u. :)

OpenStudy (anonymous):

can u check if im doing it right because i feel like i messed up somewhere?

OpenStudy (mathstudent55):

Sure. Show what you did.

OpenStudy (anonymous):

i found the slope using m= Y2-Y1/X2-X1. i got 4/7

OpenStudy (mathstudent55):

Correct, m = 4/7

OpenStudy (anonymous):

then i plugged it in y=mx+b. so now i have y=4/7x+b.

OpenStudy (anonymous):

i plugged in point (5,4) sonow i have 4=4/7x5+b

OpenStudy (anonymous):

then i got 4=20/7+b.thats were im not sure if im right.

OpenStudy (mathstudent55):

 Correct